# Sharp Analysis for KL-Regularized Contextual Bandits and RLHF

### v1

@@ Appendix D Proofs from Section 4
$\mathbb{E}_{\pi_{\theta_{*}}^{\eta}}\bigg{[}R(\theta_{*},x,a)-\frac{1}{\eta}\ln\frac{\pi_{\theta_{*}}^{\eta}(a|x)}{\pi_{0}(a|x)}\bigg{]}-\mathbb{E}_{\pi_{\widehat{\theta}}^{\eta}}\bigg{[}R(\theta_{*},x,a)-\frac{1}{\eta}\ln\frac{\pi_{\widehat{\theta}}^{\eta}(a|x)}{\pi_{0}(a|x)}\bigg{]}$
$=\frac{1}{\eta}\mathbb{E}_{\pi_{\theta_{*}}^{\eta}}\bigg{[}\ln\frac{\pi_{0}(a|x)\cdot\exp\bigl{(}\eta R(\theta_{*},x,a)\bigr{)}}{\pi_{\theta_{*}}^{\eta}(a|x)}\bigg{]}-\frac{1}{\eta}\mathbb{E}_{\pi_{\widehat{\theta}}^{\eta}}\bigg{[}\ln\frac{\pi_{0}(a|x)\cdot\exp\bigl{(}\eta R(\theta_{*},x,a)\bigr{)}}{\pi_{\widehat{\theta}}^{\eta}(a|x)}\bigg{]}$
$=\frac{1}{\eta}\mathbb{E}_{x\sim d_{0}}\big{[}\ln Z_{\theta_{*}}^{\eta}(x)\big{]}-\frac{1}{\eta}\mathbb{E}_{x\sim d_{0}}\big{[}\ln Z_{\widehat{\theta}}^{\eta}(x)\big{]}-\mathbb{E}_{x\sim d_{0}}\bigg{[}\sum_{a\in\mathcal{A}}\pi_{\widehat{\theta}}^{\eta}(a|x)\cdot\big{(}R(\theta_{*},x,a)-R(\widehat{\theta},x,a)\big{)}\bigg{]}.$
For an arbitrary reward function $f:\mathcal{X}\times\mathcal{A}\to\mathbb{R}$, let $\Delta(x,a)=f(x,a)-R(\theta_{*},x,a)$. Consider the following first derivative of $J(f)=\ln Z_{f}^{\eta}(x)-\eta\sum_{a\in\mathcal{A}}\pi_{f}^{\eta}(a|x)\cdot\Delta(x,a)$, where $Z_{f}^{\eta}(x)=\sum_{a\in\mathcal{A}}\pi_{0}(a|x)\cdot\exp(\eta\cdot f(x,a))$ and $\pi_{f}^{\eta}(a|x)\propto\pi_{0}(a|x)\cdot\exp(\eta\cdot f(x,a))$.
@@
$\quad\frac{\partial}{\partial\Delta(x,a)}\bigg{[}\ln Z_{f}^{\eta}(x)-\eta\sum_{a\in\mathcal{A}}\pi_{f}^{\eta}(a|x)\cdot\Delta(x,a)\bigg{]}$
@@
$J(R(\widehat{\theta},x,\cdot))=\ln Z_{\widehat{\theta}}^{\eta}(x)-\eta\sum_{a\in\mathcal{A}}\pi_{\widehat{\theta}}^{\eta}(a|x)\cdot\bigl{(}R(\widehat{\theta},x,a)-R(\theta_{*},x,a)\bigr{)}$
$=\ln\sum_{a\in\mathcal{A}}\pi_{0}(a|x)\cdot\exp(\eta(R(\widehat{\theta},x,a)-b(x)))-\eta\sum_{a\in\mathcal{A}}\pi_{\widehat{\theta}}^{\eta}(a|x)\cdot\bigl{(}R(\widehat{\theta},x,a)-R(\theta_{*},x,a)-b(x)\bigr{)}$
@@
From Lemma D.2, if $m\geq 32\eta^{2}D^{2}e^{B}\cdot\log(2N_{\mathcal{R}}(\epsilon_{c})/\delta)$, for any $(x,a)\in\mathcal{X}\times\mathcal{A}$ such that $\pi_{0}(a|x)>0$, it holds that
@@
Let $\epsilon_{c}=\min\{\frac{\epsilon}{2(1+c_{m,n}^{-1})e^{B}},\frac{1}{(1+c_{m,n})e^{B}\eta^{2}D^{2}}\}$. From Lemma D.4, under the condition of the theorem, with high probability the output policy $\pi_{\widehat{\theta}}^{\eta}$ is $O(\epsilon)$ optimal. ∎
@@
The proof follows the same lines as Theorem 4.4 by replacing the global-coverage condition with the local-coverage condition. It still holds that
$Q(\pi^{*})-Q(\pi_{\widehat{\theta}_{0}}^{\eta})\leq\eta\cdot\mathbb{E}_{\pi_{f}^{\eta}}\bigl{[}\bigl{(}R(\widehat{\theta}_{0},x,a)-R(\theta_{*},x,a)-b(x)\bigr{)}^{2}\bigr{]}$
@@ Appendix E Auxiliary Lemmas
Lemma E.1 (Freedman inequality).

### v2

@@ Appendix D Proofs from Section 4
$\mathbb{E}_{\pi_{\theta_{*}}^{\eta}}\bigg{[}R(\theta_{*},x,a)-\frac{1}{\eta}\log\frac{\pi_{\theta_{*}}^{\eta}(a|x)}{\pi_{0}(a|x)}\bigg{]}-\mathbb{E}_{\pi_{\widehat{\theta}}^{\eta}}\bigg{[}R(\theta_{*},x,a)-\frac{1}{\eta}\log\frac{\pi_{\widehat{\theta}}^{\eta}(a|x)}{\pi_{0}(a|x)}\bigg{]}$
$=\frac{1}{\eta}\mathbb{E}_{\pi_{\theta_{*}}^{\eta}}\bigg{[}\log\frac{\pi_{0}(a|x)\cdot\exp\bigl{(}\eta R(\theta_{*},x,a)\bigr{)}}{\pi_{\theta_{*}}^{\eta}(a|x)}\bigg{]}-\frac{1}{\eta}\mathbb{E}_{\pi_{\widehat{\theta}}^{\eta}}\bigg{[}\log\frac{\pi_{0}(a|x)\cdot\exp\bigl{(}\eta R(\theta_{*},x,a)\bigr{)}}{\pi_{\widehat{\theta}}^{\eta}(a|x)}\bigg{]}$
$=\frac{1}{\eta}\mathbb{E}_{x\sim d_{0}}\big{[}\log Z_{\theta_{*}}^{\eta}(x)\big{]}-\frac{1}{\eta}\mathbb{E}_{x\sim d_{0}}\big{[}\log Z_{\widehat{\theta}}^{\eta}(x)\big{]}-\mathbb{E}_{x\sim d_{0}}\bigg{[}\sum_{a\in\mathcal{A}}\pi_{\widehat{\theta}}^{\eta}(a|x)\cdot\big{(}R(\theta_{*},x,a)-R(\widehat{\theta},x,a)\big{)}\bigg{]}.$
For an arbitrary reward function $f:\mathcal{X}\times\mathcal{A}\to\mathbb{R}$, let $\Delta(x,a)=f(x,a)-R(\theta_{*},x,a)$. Consider the following first derivative of $J(f)=\log Z_{f}^{\eta}(x)-\eta\sum_{a\in\mathcal{A}}\pi_{f}^{\eta}(a|x)\cdot\Delta(x,a)$, where $Z_{f}^{\eta}(x)=\sum_{a\in\mathcal{A}}\pi_{0}(a|x)\cdot\exp(\eta\cdot f(x,a))$ and $\pi_{f}^{\eta}(a|x)\propto\pi_{0}(a|x)\cdot\exp(\eta\cdot f(x,a))$.
@@
$\frac{\partial}{\partial\Delta(x,a)}\bigg{[}\log Z_{f}^{\eta}(x)-\eta\sum_{a\in\mathcal{A}}\pi_{f}^{\eta}(a|x)\cdot\Delta(x,a)\bigg{]}$
@@
$J(R(\widehat{\theta},x,\cdot))=\log Z_{\widehat{\theta}}^{\eta}(x)-\eta\sum_{a\in\mathcal{A}}\pi_{\widehat{\theta}}^{\eta}(a|x)\cdot\bigl{(}R(\widehat{\theta},x,a)-R(\theta_{*},x,a)\bigr{)}$
$=\log\sum_{a\in\mathcal{A}}\pi_{0}(a|x)\cdot\exp(\eta(R(\widehat{\theta},x,a)-b(x)))-\eta\sum_{a\in\mathcal{A}}\pi_{\widehat{\theta}}^{\eta}(a|x)\cdot\bigl{(}R(\widehat{\theta},x,a)-R(\theta_{*},x,a)-b(x)\bigr{)}$
@@
By Lemma D.2, if $m\geq 32\eta^{2}D^{2}e^{B}\cdot\log(2N_{\mathcal{R}}(\epsilon_{c})/\delta)$, for any $(x,a)\in\mathcal{X}\times\mathcal{A}$ such that $\pi_{0}(a|x)>0$, it holds that
@@
Let $\epsilon_{c}=\min\{\frac{\epsilon}{2(1+c_{m,n}^{-1})e^{B}},\frac{1}{(1+c_{m,n})e^{B}\eta^{2}D^{2}}\}$. By Lemma D.4, under the condition of the theorem, with high probability the output policy $\pi_{\widehat{\theta}}^{\eta}$ is $O(\epsilon)$ optimal. ∎
@@
The proof follows the same lines as Theorem 4.7 by replacing the data coverage condition with the local-coverage condition. It still holds that
$Q(\pi^{*})-Q(\pi_{\widehat{\theta}_{0}}^{\eta})\leq\eta\cdot\mathbb{E}_{\pi_{f}^{\eta}}\bigl{[}\bigl{(}R(\widehat{\theta}_{0},x,a)-R(\theta_{*},x,a)-b(x)\bigr{)}^{2}\bigr{]},$
@@ Appendix E Auxiliary Lemmas
Lemma E.1 (Freedman’s Inequality).
@@
Lemma E.5 (Solution of KL-regularized Optimization (Proposition 7.16 of Zhang 2023)).
For any fixed $x\in\mathcal{X}$ and reward function $R$, we have
$\max_{\pi}\mathbb{E}_{a\sim\pi(\cdot|x)}\Big{[}R(x,a)-\eta^{-1}\mathrm{KL}\big{(}\pi(\cdot|x)\|\pi_{0}(\cdot|x\big{)}\Big{]}$
$=\frac{1}{\eta}\cdot\log\mathbb{E}_{a\sim\pi_{0}(\cdot|x)}\exp\big{(}\eta R(x,a)\big{)},$
where $Z_{R}(x)$ is the normalization constant and the minimizer of the loss functional is
$\pi_{R}^{\eta}(a|x)=\frac{1}{Z_{R}(x)}\pi_{0}(a|x)\exp\Big{(}\eta R(x,a)\Big{)}.$